Find three consecutive integers such that the sum of the smallest and the largest is 40.
Answer by rapaljer(4671)

Let x = smallest
x+1 = middle
x+2 = largest

x + x+2 = 40
2x+ 2 = 40
2x= 38
x=19
x+1 = 20
x+2 = 21

Check: The first and third add up to 40.

R^2